# LLM4DSR: Leveraing Large Language Model for Denoising Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation systems fundamentally rely on users’ historical interaction sequences, which are often contaminated by noisy interactions. Identifying these noisy interactions accurately without additional information is particularly difficult due to the lack of explicit supervisory signals to denote noise. Large Language Models (LLMs), equipped with extensive open knowledge and semantic reasoning abilities, present a promising avenue to bridge this information gap. However, employing LLMs for denoising in sequential recommendation introduces notable challenges: 1) Direct application of pretrained LLMs may not be competent for the denoising task, frequently generating nonsensical responses; 2) Even after fine-tuning, the reliability of LLM outputs remains questionable, especially given the complexity of the task and th inherent hallucinatory issue of LLMs.
To tackle these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendation using LLMs. We constructed a self-supervised fine-tuning task to activate LLMs’ capabilities to identify noisy items and suggest replacements. Furthermore, we developed an uncertainty estimation module that ensures only high-confidence responses are utilized for sequence corrections. Remarkably, LLM4DSR is model-agnostic, allowing the corrected sequences to be flexibly applied across various recommendation models. Extensive experiments validate the superiority of LLM4DSR over existing methods across three datasets and three recommendation backbones.
@@ 1. Introduction
Given the powerful capabilities of LLMs, this work explores a innovative question: Can LLMs effectively function as denoisers for sequential recommendation? Sequential recommendation critically relies on the accuracy of users’ historical interaction sequences to predict subsequent items. However, in practice, these sequences often contain noise due to various factors — e.g., being attracted by clickbait (Wang et al., 2021b), influenced by item prominent positions (Chen et al., 2023), or from accidental interactions (Lin et al., 2023). Such noise can significantly mislead recommendation models and degrade their performance, thus underscoring the necessity for effectively denoising.
@@
The task of denoising is inherently difficult. A primary issue is the absence of labels for noisy data, resulting in a lack of clear signals regarding the nature of the noise. Recent approaches to denoising in sequential recommendation have either relied on human-designed heuristics (Zhang et al., 2024a, 2022; Sun et al., 2021), which require extensive expertise and often lack precision; or have been integrated with recommendation models guided by the recommendation objective (Zhang et al., 2022; Yuan et al., 2021; Chen et al., 2022; Lin et al., 2023), which may still be heavily influenced by the noisy data, particularly since the labels of the recommendation objective (i.e., the next item) might also be contaminated. Thus, this work explores a new avenue of leveraging LLMs for recommendation denoising, which can potentially compensate for this knowledge gap. LLMs, with their extensive open knowledge and advanced semantic reasoning abilities (Brown et al., 2020), are equipped to identify abnormal interactions from a semantic perspective. For example, if a user’s viewing history primarily consists of comedy films, an occasional interaction with a horror film might be attributable to an accidental click. LLMs can detect such anomalies based on their understanding of item semantics and common knowledge. Moreover, by leveraging the generative capabilities of LLMs, they can not only identify noisy items but also potentially suggest replacements that align more closely with user preferences.
@@
However, employing pre-trained LLMs for sequence denoising also presents significant challenges:
@@
Challenge 1: Direct application of pretrained LLMs may not be competent for denoising tasks. The significant gap between the pre-training objectives of LLMs and the specific requirements of sequence recommendation denoising can make their direct application problematic. Initial tests with a pre-trained Llama3-8B model (Dubey et al., 2024) indicated that a large portion (40%) of the responses were nonsensical. Figure 1 illustrates examples where LLMs typically misidentify all items in a sequence as noise.
@@
Challenge 2: Responses from LLMs might not always be reliable.. Even with fine-tuning, LLMs are susceptible to the notorious phenomenon of hallucination (Ji et al., 2023), where they might incorrectly label non-existent items as noise or suggest non-existent replacements. Moreover, given the complexity of the recommendation denoising task, the accuracy of an LLM-based denoiser still has an upper-bound. Blindly trusting LLMs for corrections may not be advisable.
@@
To address these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendations using LLMs. To overcome the first challenge, we adapt pre-trained LLMs to the denoising task through a self-supervised fine-tuning process. We construct an instruction dataset by replacing a proportion of items in sequences with randomly selected alternatives, which are then used for fine-tuning LLMs to identify noisy items and suggest appropriate replacements. This approach activates the potential denoising ability of LLMs. For the second challenge, we introduce an uncertainty estimation module to assess the reliability of the identified noisy items, ensuring that only high-confidence judgments are used to correct sequences for subsequent recommendations. Remarkably, LLM4DSR is model-agnostic, making the corrected sequences flexibly applicable across various recommendation models. Our extensive experiments demonstrate that LLM4DSR significantly enhances the performance of both traditional sequential recommendation models and advanced LLM-based models.
@@ 2. Preliminary
In this section, we introduce key definitions in this work.
@@
Sequential Recommendation Let be the set of items. We use to denote -th historical temporal interaction sequence where is the -th item interacted within sequence , and represents the sequence’s length. We denote a subsequence of as . For each sequence , sequential recommendation takes sequence data as input and outputs the item that the user is likely to interact with next, which is expected as . For the sake of brevity, we will use to replace in the following text.
Denoise Task Formulation For a given sequence , the objective of the denoising task is to identify and remove the noise within the sequence and ultimately generate a noiseless sequence . Previous explicit denoising tasks simply performed deletion operations on the sequence, resulting in a subsequence that is shorter than the original sequence, i.e., and .
In this paper, the denoising task is accomplished by substituting the identified noise with a more appropriate item. The reason is it can effectively reduce the impact of erroneous deletions, and the extra items introduced by the replacement operation may indeed introduce some useful information benefiting sequence recommendation. We also empirically validate its effectiveness in our experiments. The replacement operation maintains the sequence length unchanged, i.e., and modify the original sequence by
@@
where indicates whether each item at a given position is noise. indicates that the denoising model classify the item as noise, while indicates not. denotes the replacement item generated by the model for .
3. Method
In this section, we outline the methodology of LLM4DSR. Section 3.1 details the training process for the LLM to enable its denoising capabilities. In Section 3.2, we present the uncertainty estimation module, which is designed to enhance the accuracy of the denoising process. Finally, Section 3.3 describes the procedure for replacing noise samples with the suggested items provides by LLMs.
3.1. Training LLM as a Denoiser
Due to the significant discrepancy between the original training objectives in pre-training phase and denoising tasks, LLMs inherently lack the capability for denoising. Previous studies have indicated that fine-tuning with a small amount of data can enable LLMs to quickly adapt to new tasks (Brown et al., 2020). In this section, we design a self-supervised task and employ supervised fine-tuning (SFT) to empower LLMs with denoising ability.
For an original sequence in the training set, we randomly select an element and replace it with an item randomly sampled from the item set , generating a modified sequence . The training objective of the LLM is to identify the replaced item given the modified sequence and predict the original item at that position. This self-supervised task can be expressed in natural language as follows:
@@ 3.1. Training LLM as a Denoiser
Here, represents the title of item . We denote the token sequence of input part as and output part as . The token sequences for ”Noise Items:” and ”Suggested Items:” in the output are defined as and , respectively. Then we can fine-tune the LLMs on the aforementioned corpus and the objective function can be formalized as follows:
@@
Training on the aforementioned self-supervised tasks empowers the LLMs to effectively understand and perform denoising tasks. The prompt ”The user has interacted with the following items before:” provides the model with information about the user’s historical interactions. ”Noise Items:” guides the model to identify noise within the sequence, as these noise items typically deviate significantly from other items. ”Suggested Items:” encourages the model to infer the user’s interests based on the most items in the sequence, thereby suggesting items that align more closely with the user’s preferences.
@@
While the LLM has developed the capability to denoise using the ”find and replace” task, one limitation is that it can only address a single noise item per sequence. This is because the training data used during the fine-tuning phase contains only one item in the output, leading the model to strictly output one item during the inference phase as well 111We also attempted to perform multiple replacements on a sequence in the self supervised task to enable the LLM to output multiple noise items, but we found that this approach yielded very low accuracy.. Consequently, the model is unable to effectively manage situations where multiple noise items are present. Furthermore, due to the hallucination phenomenon inherent in LLMs, directly relying on the model’s output to identify noise is unreliable. The model may generate fake items that are not actually present in the sequence, or it may lack high confidence in the given item, thereby compromising the accuracy of the denoising results. To mitigate these issues, we have developed an uncertainty estimation module, which calculates the probability of each item in the sequence being noise and use them to identify the noise.
3.2. Uncertainty Estimation
Relying exclusively on responses from LLMs for denoising is both imprecise and inflexible. We developed an uncertainty estimation module designed to gauge the model’s confidence in its response.
For a given sequence , we define the probability of item being noise as . Considering that LLMs are capable of modeling text probabilities, we utilize the LLM to provide the probability of appearing after the prompt ”Noise Items:” in the context of a user-provided sequence . This probability serves as an indicator of the LLM’s confidence in identifying the item as noise and can be treated as an estimation of . Formally,
@@ 3.2. Uncertainty Estimation
where denotes the token vector of the item and is the total length of .
The calculation of necessitates a forward pass for each individual item in the sequence, which can be computationally expensive for LLMs. Based on the observation that the first token of different items often varies for most sequences, we can simplify the probability computation. For such sequences, once the first token is given, the subsequent tokens are fully determined. This is because the generated noise item must be included among the sequence items. Formally,
Thus we can get,
Given the aforementioned simplification, we only need to perform a single forward pass using the prompt to obtain the probability values of all items in the sequence being noise, no need for performing one forward pass for each item in the sequence individually.
By modeling text probabilities using a LLM, we are able to determine the likelihood of each item in a sequence being noise. Ideally, a higher probability indicates a greater likelihood of the item being noise. We validated this hypothesis through experiments conducted on a dataset with artificial noise (noise label are available), the details of datasets can be found in Section 4.1.1. As shown in Figure 3, both the number and proportion of noise samples increase as the probability values increase, while the number of clean samples decreases. This experiment confirms the reliability of the probability provided by the LLM.
We propose setting a threshold to classify items with probability values exceeding this threshold as noise. Formally, for a sequence , we denote the probability values for each item as . We introduce a hyperparameter as the threshold and classify items with a probability greater than as noise, i.e.,
Uncertainty estimation of leveraging model-generated probability, rather than relying solely on response-based discrimination, enhances the accuracy of noise detection while also increasing the algorithm’s flexibility. In certain sequences, the model may exhibit high confidence that multiple items are noise, resulting in several items being modified, whereas in others, the confidence levels across all items may be relatively low, leading to no modifications being made.

### v2

@@ Abstract.
Sequential Recommenders generate recommendations based on users’ historical interaction sequences. However, in practice, these collected sequences are often contaminated by noisy interactions, which significantly impairs recommendation performance. Accurately identifying such noisy interactions without additional information is particularly challenging due to the absence of explicit supervisory signals indicating noise. Large Language Models (LLMs), equipped with extensive open knowledge and semantic reasoning abilities, offer a promising avenue to bridge this information gap. However, employing LLMs for denoising in sequential recommendation presents notable challenges: 1) Direct application of pretrained LLMs may not be competent for the denoising task, frequently generating nonsensical responses; 2) Even after fine-tuning, the reliability of LLM outputs remains questionable, especially given the complexity of the denoising task and the inherent hallucinatory issue of LLMs.
To tackle these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendation using LLMs. We constructed a self-supervised fine-tuning task to activate LLMs’ capabilities to identify noisy items and suggest replacements. Furthermore, we developed an uncertainty estimation module that ensures only high-confidence responses are utilized for sequence corrections. Remarkably, LLM4DSR is model-agnostic, allowing corrected sequences to be flexibly applied across various recommendation models. Extensive experiments validate the superiority of LLM4DSR over existing methods.
@@ 1. Introduction
Given the powerful capabilities of LLMs, this work explores a innovative question: Can LLMs effectively function as denoisers for sequential recommendation? Sequential recommendation critically relies on the accuracy of users’ historical interaction sequences to predict subsequent items. However, in practice, these sequences are often contaminated by noise noise due to various factors — e.g., being attracted by clickbait (Wang et al., 2021b), influenced by item prominent positions (Chen et al., 2023), or from accidental interactions (Lin et al., 2023). Such noise can significantly mislead recommendation models and degrade their performance. Figure 2 provides empirical evidence of this impact: with only a small proportion (e.g., 20%) of randomly selected noisy items, we observe performance drops of 41% and 36% in the sequential recommendation models SASRec and LLaRA, respectively. The importance of denoising these sequences cannot be overstated.
@@
Despite its importance, the task of denoising is inherently challenging. A primary issue is the absence of labels for noisy data, resulting in a lack of clear signals regarding the nature of the noise. Recent approaches to denoising in sequential recommendation either rely on on human-designed heuristics (Zhang et al., 2024a, 2022; Sun et al., 2021), which require extensive expertise and often lack precision; or are trained jointly with recommendation models guided by the recommendation objective (Zhang et al., 2022; Yuan et al., 2021; Chen et al., 2022; Lin et al., 2023), which may still be heavily influenced by the noisy data, particularly since the labels of the recommendation objective (i.e., the next item) might also be contaminated. Thus, this work explores a new avenue of leveraging LLMs for recommendation denoising, which can potentially compensate for this knowledge gap. LLMs, with their extensive open knowledge and advanced semantic reasoning abilities (Brown et al., 2020), are equipped to identify abnormal interactions from a semantic perspective. For example, if a user’s viewing history primarily consists of comedy films, an occasional interaction with a horror film might be attributable to an accidental click. LLMs can detect such anomalies based on their understanding of item semantics and common-sense knowledge. Moreover, by leveraging the generative capabilities of LLMs, they can not only identify noisy items but also potentially suggest replacements that align more closely with user preferences.
@@
However, employing pre-trained LLMs for denoising sequential recommendation also presents significant challenges:
@@
Challenge 1: Direct application of pretrained LLMs may not be competent for denoising tasks. A common approach to utilizing LLMs involves structuring the denoising task as a prompt and directly feeding it into the LLMs. However, the significant gap between the pre-training objectives of LLMs and the specific requirements of sequence recommendation denoising can make such direct application problematic. In our preliminary experiments with the pretrained Llama3-8B model (Dubey et al., 2024), we observed that a significant portion (over 40%) of the model’s responses were nonsensical. Figure 1 presents typical failure cases, where the LLM frequently misclassifies all items in a sequence as noise.
@@
Challenge 2: Responses from LLMs might not always be reliable. Even with fine-tuning, LLMs are susceptible to the notorious phenomenon of hallucination (Ji et al., 2023), where they might incorrectly label non-existent items as noise or suggest non-existent replacements. Moreover, given the complexity of the recommendation denoising task, the accuracy of an LLM-based denoiser still has an upper-bound. Blindly trusting LLMs for corrections may not be advisable.
@@
To address these challenges, we propose LLM4DSR, a tailored approach for denoising sequential recommendations using LLMs. To overcome the first challenge, we adapt pre-trained LLMs to the denoising task through a self-supervised fine-tuning process. We construct an instruction dataset by replacing a proportion of items in sequences with randomly selected alternatives, which are then used for fine-tuning LLMs to identify noisy items and suggest appropriate replacements. This approach activates the potential denoising ability of LLMs. To tackle the second challenge, we introduce an uncertainty estimation module to assess the reliability of the identified noisy items, ensuring that only high-confidence judgments are used to correct sequences for subsequent recommendations. Notably, LLM4DSR is data-centric and model-agnostic, which can be considered a data preprocessing technique. The corrected sequences can be seamlessly applied across various recommendation models. Our extensive experiments demonstrate that LLM4DSR significantly enhances the performance of both traditional sequential recommendation models and advanced LLM-based models.
@@
The remainder of this article is organized as follows: Section 2 outlines the basic definitions of the tasks and methods involved. The proposed method, LLM4DSR, is introduced in Section 3. Section 4 presents our experimental results and subsequent discussions. Related work is reviewed in Section 5. Finally, we conclude the article and suggest future research directions in Section 6.
| Notations | Descriptions |
| the -th item in the item set | |
| the historical temporal interaction sequence | |
| the noiseless sequence of generated by the model | |
| the modified sequence of generated by randomly replacing | |
| the -th item in the sequence | |
| whether the interaction is noise predicted by the model | |
| the replacement item generated by the model for | |
| the textual description of the item | |
| the token sequence of the input part of the prompt | |
| the token sequence of the out part of the prompt | |
| the token sequence of the token sequence of the noise item given by the LLM | |
| the model parameter | |
| the token sequence of the textual description of the item | |
| the probability threshold to classify items as noise | |
| noise ratio of artificially noisy dataset | |
| the maximum number of items modified in a single sequence | |
| the ranking position of the item for the user . |
@@ 2. Preliminary
In this section, we present the background of the task of sequential recommendation denoising and the large language models. Tabel LABEL:notation summarizes the notations involved in this article.
2.1. Background of Sequential Recommendation
@@
Our problem definitions directly refer to recent work (Kang and McAuley, 2018; Sun et al., 2019; Bao et al., 2023a). Given a sequential recommender system, let be the set of items. We use to denote -th historical collected interaction sequence where denotes the -th item interacted within sequence , and represents the sequence’s length. We denote a subsequence of as . For each sequence , sequential recommendation takes sequence data as input and outputs the item that the user is likely to interact with next, which is expected as . For the sake of brevity, we will use to replace in the reminder of the paper.
2.2. Sequential Recommendation Denoising.
The presence of noisy interactions in user behavior sequences can impact model performance. As shown in Figure 2, both traditional sequential recommendation models and recent large language model-based recommendation models exhibit a significant decline in performance as the noise ratio increases. Therefore, developing denoising techniques for sequential recommendation is an important task.
Recent work on sequential recommendation denoising can be mainly categorized into two types: 1) the explicit denoising and 2) the implicit denoising. For a given sequence , the objective of the explicit denoising strategies (Sun et al., 2021; Zhang et al., 2022; Lin et al., 2023) target at identifying and removing the noise within the sequence and ultimately generate a noiseless sequence . Implicit denoising methods (Yuan et al., 2021; Chen et al., 2022; Zhou et al., 2022; Han et al., 2024) do not directly remove items from the sequence. Instead, they reduce the impact of noise with respect to the final representation by decreasing the weight of noise items, or by filtering noise at the representation level through filters.
Traditional denoising methods encounter significant challenges due to the absence of labeled noise data. These methods often rely on heuristic prior knowledge for noise identification, which requires extensive expertise and frequently lacks precision. Alternatively, some denoising approaches employ recommendation loss as a supervisory signal during the denoising process; however, this can be adversely affected by the presence of noise. Moreover, explicit denoising methods risk information loss from sequences due to potential erroneous deletions, which can degrade model performance.
To address these limitations, this work explores leveraging LLMs in sequential recommendation denoising. The advantages of this strategy are multifaceted: 1) LLMs possess extensive open knowledge and advanced semantic reasoning abilities, which can potentially bridge the information gap in the denoising task; 2) The denoising task can be decoupled from the learning of recommendation models, serving as a data preprocessing step, thus allowing the corrected sequence to benefit various sequential recommendation models; 3) Owing to the generative ability of LLMs, the method can not only identify noisy items but also potentially suggest replacements. This can effectively reduce the impact of erroneous deletions, and the additional items introduced by the replacement operation may indeed provide useful information, benefiting sequence recommendation. We also empirically validate the effectiveness of such generation in our experiments.
Formally, let denote the corrected sequence of the original sequence , where each represents the -th item of the corrected sequence. The denoising process in this paper can be expressed as follows:
@@
where indicates that the denoising model classify the item as noise, while indicates not. denotes the replacement item generated by the LLMs for .
2.3. Background of Large Language Model
Large Language Models (LLMs) have become transformative tools in natural language processing. These models, built on the transformer architecture (Vaswani, 2017), effectively capture relationships in sequential data and can easily scaled to unprecedented sizes, reaching billions of parameters (Dubey et al., 2024; Achiam et al., 2023). As the number of parameters increases, LLMs exhibit emergent abilities not present in traditional NLP models, leading to significant improvements in language understanding and generation tasks (Wei et al., 2022a). Trained on vast corpora of text from diverse sources, LLMs are able to capture a wide range of linguistic patterns and contextual nuances. They have demonstrated remarkable capabilities in tasks such as translation (Zhu et al., 2023), summarization (Basyal and Sanghvi, 2023), and question-answering (Huo et al., 2023), making them a pivotal component of modern AI research and applications.
In addition, LLMs pre-trained on extensive datasets exhibit strong zero-shot or few-shot capabilities for downstream tasks, requiring only minimal fine-tuning to quickly adapt to specific tasks (Zhang et al., 2023a). Instruction tuning (Chung et al., 2024) is a fine-tuning method that enables LLMs to better understand and follow human instructions. By training the model on a labeled dataset of instructional prompts and corresponding outputs, instruction tuning significantly improves the model’s performance in specified instruction formats. This approach not only refines the model’s outputs but also enhances its interpretability and reliability in practical applications. Specifically, the instruction data is organized in the format , where represents the text-based instruction and represents the corresponding response. LLMs have demonstrated a significant improvement in performance on specific tasks when trained with a limited amount of instructional data (Zhang et al., 2023a).
3. Methodology
In this section, we present the proposed LLM4DSR. We first employ a self-supervised instruction tuning approach to enable the denoising capability of the LLM (Subsection 3.1), and then introduce an uncertainty estimation module to improve the reliability of the denoising process (Subsection 3.2). Finally, we describe the procedure for replacing noisy samples with items suggested by the LLM (Subsection 3.3).
3.1. Fine-tuning LLM as a Denoiser
To better activate the denoising capabilities of LLMs, we propose leveraging an instruction tuning strategy. Instruction tuning has been widely adopted in various LLM-based applications, and has proven effective in enabling LLMs to quickly adapt to new tasks (Zhang et al., 2023a). Specifically, the template for the instruction data in this task can be formulated as follows:
where represents the textual description of item (e.g., titles). Here, we organize the language descriptions of sequences as instruction inputs and expect the LLMs to identify noisy items and suggest replacements. Specifically, the prompt ”The user has interacted with the following items before:” provides the model with information about the user’s historical interactions. ”Noise Items:” guides the model to identify noise within the sequence from the semantic context. ”Suggested Items:” encourages the model to infer the user’s interests based on the sequence and suggest replacements that align more closely with the user’s preferences.
However, one challenge of such an instruction fine-tuning strategy is the lack of explicit signals denoting the ground truth of the noise items and ideal replacement items. To address this, we propose leveraging a self-supervised strategy to construct instruction data. We randomly select an item from the original sequences and replace it with a noise item that is randomly sampled from . LLMs can be fine-tuned on the corrupted sequence, aiming to predict the modified noise item and make predictions of the original item. The schematic diagram of the self-supervised task is shown in Figure 3. In this way, LLMs can effectively understand the task of sequential recommendation denoising and generalize to identify potential abnormal items in the original sequence and make corrections, which would significantly boost recommendation performance.
@@ 3.1. Fine-tuning LLM as a Denoiser
Formally, let denote the token sequence of input part as and output part as . Then we can fine-tune the LLMs on the aforementioned corpus and the objective function can be formalized as follows:
@@
3.2. Uncertainty Estimation Module
@@
While the aforementioned instruction tuning has activated the denoising potential of LLMs, they still face two challenges during applications: 1) Given the notorious hallucination phenomenon inherent in LLMs and the complexity of the denoising task, directly relying on the model’s response to identify noise items might not be reliable or precise. The model may generate fake items that are not actually present in the sequence, or it may lack high confidence in the given item, thereby compromising the accuracy of the denoising results. 2) The response is limited to addressing a single noisy item per sequence. This is because the training data used during fine-tuning contains only one noisy item in the output, leading the model to consistently output a single item during inference 111We also have attempted to construct instruction data with multiple noisy items and guide the LLM to identify several noise items. However, this approach resulted in very low accuracy. We hypothesize that this is due to the increased difficulty of directly identifying a varying number of noisy items.. Consequently, this strategy is inflexible in handling cases where multiple noisy items are present.
To tackle these issues, we introduce an uncertainty estimation strategy. Rather than directly utilizing the response from LLMs, which could be imprecise and inflexible, we opt to scrutinize the probability of each item in the sequence being regarded as a noise item by LLMs:
@@
where we examine the generative probability of the description of the item . Here, represents the token sequence of the noise item. denotes the token sequence of , denotes the total length of and represents the -th token of .
Considering the calculation of can be time-consuming due to the need to iterate over each token, in practice, we may simplify the process by calculating the probability of only the first token. This is because the first token of the items in a sequence is typically unique, suggesting that the generated item description can be determined primarily by the first token. Our empirical analysis also demonstrates that the generative probability for the remaining tokens is approximately 1.
The probability reflects the confidence of the LLM’s judgment — the higher the probability, the more likely the item is to be a noisy item. To empirically validate its effectiveness, we conducted experiments on a dataset with an artificial noise ratio of 10% 222We use a dataset with artificially injected noise because only in this setting the ground-truth labels for noisy items are available.. The details of the dataset are provided in Section 4.1.1. As shown in Figure 5, as the probability increases, both the number and proportion of noise samples increase, while the number of clean samples decreases.
Thus, in our approach, is used to identify noisy items within the sequence. Formally, we apply a threshold for classification:
This uncertainty-based strategy mitigates the hallucination issue inherent in LLMs by avoiding reliance on potentially erroneous generative outputs. Besides, this strategy provides flexibility in handling multiple noisy items, as the model can assign a high to multiple items in the sequence.